# “Add more config detail”: A Taxonomy of Installation Instruction Changes

### v1

@@ 3.2 README Commit Classification
Based on the information obtained from the headers, we developed semi-automated labelling heuristics implemented as a program. Starting from the $\langle h1\rangle$ headers, this program recursively examines whether the modified sections in the README commit are in the “relevant” set or not. If any of the modified sections has a “relevant” header, the program will mark the commit as relevant. Conversely, if all modified sections are categorised under the “irrelevant” headers, the commit is marked accordingly as irrelevant. Otherwise, the program cannot determine and proceeds to the next level of headers, repeating the decision process. If it still cannot ascertain the relevance after inspecting the headers at level 4, it assigns a “not sure” label, indicating that human annotation is required.
@@
To verify the accuracy of this program, we manually annotated four repositories, and the program reached an accuracy of 92.4% (133/144). Subsequently, we applied this program to all repositories, resulting in 7,161 relevant, 8,583 irrelevant and 9,465 “not sure” commits generated by the semi-automatic labelling heuristics.
@@
To annotate the data points categorised as “not sure”, the first author randomly selected 50 records and three authors independently performed the annotation. The outcome of the annotation reached a Fleiss’ Kappa score of 0.69, indicating a good level of agreement. Three authors further discussed the disagreed entries and came to an eventual agreement. Subsequently, the first author continued to annotate the remaining commits. Finally, within the dataset of 25,230 README commits, 8,673 commits were identified as relevant and 16,557 as irrelevant. The high accuracy rate (over 92.4%) and high recall (82.6%) indicate that our heuristic is quite capable of categorising the README documentation.
3.3 Dataset Characteristics
@@
Before the details of data analysis, we briefly introduce the characteristics of our curated dataset. The 8,673 identified relevant commits from the previous step cover 378 of 400 sampled repositories, that is, 94.5% of the repositories contain at least one README commit modifying installation-related sections. On average, each repository contains 23 relevant README commits, with a standard deviation of 25.7, indicating a skewed distribution. Specifically, with a minimum of 1 and a maximum of 224, the 25th, 50th, and 75th percentiles are 8, 14 and 31 commits, respectively.
@@
Therefore, we performed a cluster sampling [41], splitting the dataset into four clusters: repositories with README commits from the minimum to the first quantile, from the first quantile to the second quantile, from the second quantile to the third quantile, and from the third quantile to the maximum. Within each cluster, we performed a random sample with the sample size required to achieve a 95% confidence level (confidence interval: 5), resulting in samples of 206, 280, 324, and 358 README commits for each respective cluster. In total, we conducted a large-scale qualitative study on 1,168 README commits.
We calculated the correlation coefficient between the number of relevant README commits and the number of stars, as well as with the number of years the project has existed. The Pearson correlation coefficient reached 0.19 and 0.01, respectively. This statistic indicates that projects that actively participate in documentation activities, particularly for the installation-related sections, have a positive correlation to the project’s popularity, while the number of relevant README commits does not necessarily grow with the years of the project.
@@ 3.4 Data Analysis
We adopted thematic analysis [9] for qualitative analysis of our data, as it aligned well with our goal of deriving themes for update events. The first author conducted the data analysis and created spreadsheets containing all raw data and codes. The spreadsheets were shared among all authors, and the second and third authors cross-validated all codes, concepts, and categories to reduce bias and increase the reliability of the findings [42]. Concepts and categories were discussed among authors through regular meetings and finalised after several rounds of revisions.
@@
We started our analysis by searching each commit on GitHub and recording the text summarising the changes from the perspective according to our research question: the updates for the installation-related sections. We call this process “key points extraction and getting familiar with the data”. Then we rigorously performed open coding to assign each README commit with one to many succinct descriptions of their update modifications. This process took into account both the key points that were extracted and the content of the README commits themselves. The additional investigation into the raw README commit data is to address any unclear key points and to double-check whether any part of the modification is missing.
During the coding procedure, we constantly compared the emerging codes within and across different README commits to synthesise a higher level of abstraction called concepts, and later the highest level of abstract called categories. Figure 4(a) illustrates examples of the procedure of generating one of the categories called “Pre-installation Instruction Updates”. Figures 4(b) and Figure 4(c) further explain the process for how we assign codes for different changes within the commit. We have made our analysed data available. Please refer to Replication Package section for more details.
@@
4 Results
Table II presents the hierarchical taxonomy of 189 codes that we identified from the 1,168 data points to answer our research question. The updates are categorised into six main groups: (1) pre-installation instruction; (2) installation-related instruction; (3) post-installation instruction; (4) help information; (5) document presentation; and (6) external resource management.
To give readers a better understanding of our analysed result, Figure 6 synthesises our analysis in a timeline fashion, corresponding to pre-installation, installation, and post-installation, where different software development tasks are included in each phase. Moreover, various information perspectives regarding the README files are also displayed within this figure, including instruction, help information, external resource links, and document presentation.
For each of the categories, we describe representative examples and discuss implications for practitioners and researchers.
It is worth noting that, compared to other qualitative studies [5, 6, 1, 12], our work focuses on modifications made to Markdown-formatted text, which involves two pieces of documents. Due to the substantial space required to illustrate examples of README updates, we will only include figures for few examples. Additional examples are made accessible through URLs that we have been formatted as clickable links in Table II. We have also included the URLs and their corresponding screenshots in our replication package.
@@ 4 Results
Category Concept # Codes Frequency Example ID
Pre-installation instruction Dependency Setup 23 527 1, 2, 3, 4
Platform Setup 11 82 5, 6, 7
Virtual Machine and Hardware 6 19 ——
Total 40 628
Installation-related instruction Source code 23 267 8, 9
Published package 14 162 10, 11, 12, 13
Total 37 429
Post-installation instruction Project running instruction 12 185 14 15
Installation check 3 13 16
Running Tests 6 32 ——
Deployment Instruction 3 9 ——
Upgrade Instruction 4 6 ——
Uninstall Instruction 2 3 ——
Total 30 248
Help information Tutorial updates 6 70 17
Explanation 14 125 18, 19
Notes and troubleshooting 16 70 20, 21
Total 36 265
Document Presentation Documentation structural modification 7 197 22, 23
Content formatting and cleaning 10 119 24, 25, 26, 27
Error fixing 2 61 28, 29
Text rephrasing and editing 5 70 30, 31
Documentation debt 4 30 32
Total 28 477
External resources management Project external documentation 3 120 33
Project non-documentation artefact 11 136 34
Third-party external resources 4 110 35
Total 18 366
4.1 Pre-installation Instruction
@@
Dependency Setup (527) [^4] Number in the parenthesis represents the frequency of this concept.. Dependencies enable code reuse and require maintenance effort as the project and other dependencies evolve [43]. In our analysed dataset, more than 40% of the documentation updates contain dependency-related instructional modifications, indicating the importance of this perspective.
@@
Apart from the programming language, other dependencies setup include external libraries and modules (e.g., API), as well as third-party software (e.g., database) that projects rely on. The inclusion of new features, resolution of technical debt during project development, and the updates of constantly evolving dependencies collectively contribute to the volatility of the dependency management information within the README files. Indeed, most updates in this aspect involve updating dependency information, as indicated in Example 1, and instructions on installation of dependencies as indicated in Example 2. We observed a common practice of manifesting dependencies by listing them (e.g., in forms of plain texts, bullet points, or tables), with addition, updates, and deletion to reflect the corresponding changes. Meanwhile, the patterns for providing the installation instruction of dependencies are quite diverse. Centralised commands (e.g., requirements.txt) or separate installation commands for each dependency are widely used. However, maintaining the separate command style requires more effort, as individual attention to each dependency is needed, making modifications potentially more time-consuming and the process more complicated.
In Example 3, the commit adds configuration instructions for the database on which the project relies. This underscores the importance of incorporating essential configuration details when dependencies are not ready for use following the dependency installation step, which is a common scenario for third-party software dependencies. Interestingly, some examples, such as commit Example 4, add guidelines to verify the success of dependency installation. This type of instruction becomes more valuable for critical dependencies, especially when the installation and configuration instructions become intricate.
@@
Due to the diverse source of OSS project participants, encountering different local environments with different installed packages becomes inevitable, rendering the project installation and dependency management a challenging issue. In our analysed data, we found that 48 instances involve modification in setting up virtual environment, and a majority of them are related to newly added instructions such as the one indicated in Example 5. Furthermore, Example 6 explicitly mentions the virtual environment as their recommended way of installation of the project.
Operating system information is another platform setup concern. As project users might come from a diverse distribution of operating systems, this information becomes a critical prerequisite before project installation and usage. We observed efforts that mostly focused on adding operating system compatibility information (e.g., BiglyBT comes in several editions for different operating systems. Mac OSX, Linux and Windows use the full BiglyBT-API based on Java 8.) or update operating system version information as indicated in Example 7, which updated the support operating system from Debian 9 and 10 to Debian 10 and 11.
@@
4.2 Installation-related Instruction
@@
Installation-related Instructions from Source Code Operations (267). Installation from source code requires users to obtain the source code and then locally execute a set of predefined operations, encompassing compilation, building, and other procedures, to install the software project. While projects may use various technologies, including Docker, Pip, and Maven, we identified in our analysed data that source code-level installation is the most commonly adopted method which the projects tend to initially provide. In addition, the inclusion of newly added instructions also accounts for various operating systems. Given that commands for Windows systems typically differ from those for Linux, additional information in the README file instructions or even the source code is needed to facilitate a smooth installation experience for different users.
Documentation efforts are also observed to provide better encapsulated installation alternatives that are easier to install, as well as versions of project targeting for different user groups. For example, in commit Example 8, the installation option initially provided includes obtaining the source code and performing compilation. The newly added instruction introduces an alternative approach specifically for building and running a Docker image from the Dockerfile within the source code. This provides an integrated installation experience, alleviating concerns related to compilation and other dependency management issues.
Regarding the update behaviours for the installation instructions from source code, a wide range of patterns are identified, which are more project-specific. For example, modifications within code instructions covering project name, file path for executables, and configuration file names are usually trivial, reflecting changes at the file level and repository level. A simple example in commit Example 9 migrated the codebase from GitLab to GitHub, while also changing the project name from atoml-local to catlearn, leading to the corresponding adjustment of the commands in the README file. Updates to the source code installation option also involve adjustments of parameters, with various patterns observed for this type of change. As an illustrative example, commit Example 9 altered the Docker compose command with an additional “-d” command indicating a detached mode to run the Docker container.
@@
Regarding the update operations within the published package installation instructions, we identified a predominant pattern which focuses on updating the project version. Trivial updates are usually found in this type of modification, with merely a version number increment for repository’s latest stable version, and sometime combined with development version. For example, both the stable version and the snapshot version were incremented within the code instruction for the Maven dependency of the repository in this commit Example 11.
Apart from the version modification, other noticeable update operations for published package installation include changes for package names or package source and changes for package management technology instructions. For example, the path to the image on DockerHub was updated in commit Example 12, reflecting the changes in the source of the project. As an example of the change in technology instruction, in commit Example 13, the Gradle command compile is updated to implementation as the previous keyword was deprecated. The rest of the updates are more specific to reflect the project development status.
Summary of installation-related instruction results: The concept of installation-related instruction from source code operation is the most commonly-adopted choice in our analysed data. Their update behaviours varies, covering factors including project meta-data changes and parameter changes. On the other hand, the concept of installation-related instructions from published package operation is less frequent, and they are usually used as the additional operation on top of the source code installation alternative.
4.3 Post-installation Instruction

### v3

@@ III-B README Commit Classification
Based on the information obtained from the headers, we developed a semi-automated labelling heuristic. Starting from the $\langle h1\rangle$ headers, this program recursively examines whether the modified sections in the README commit are in the “relevant” set or not. If any of the modified sections has a “relevant” header, the program will mark the commit as relevant. Conversely, if all modified sections are categorised under the “irrelevant” headers, the commit is marked accordingly as irrelevant. Otherwise, the program cannot determine and proceeds to the next level of headers, repeating the decision process. If it still cannot ascertain the relevance after inspecting the headers at level 4, it assigns a “not sure” label, indicating that human annotation is required.
@@
To verify the accuracy of this program, we manually annotated four repositories, and the program reached an accuracy of 92.4% (133/144). Subsequently, we applied this program to all repositories, resulting in 7,161 relevant, 8,583 irrelevant and 9,465 “not sure” commits generated by the semi-automatic labelling heuristic.
@@
To annotate the data points categorised as “not sure”, the first author randomly selected 50 records and three authors independently performed the annotation. The outcome of the annotation reached a Fleiss’ Kappa score of 0.69, indicating a good level of agreement. The three authors further discussed the disagreed entries and came to an eventual agreement. Subsequently, the first author continued to annotate the remaining commits. Finally, within the dataset of 25,230 README commits, 8,673 commits were identified as relevant and 16,557 as irrelevant. The high accuracy rate (over 92.4%) and high recall (82.6%) indicate that our heuristic is quite capable of categorising the README documentation.
III-C Dataset Characteristics
Before the details of the data analysis, we briefly introduce the characteristics of our curated dataset. The 8,673 identified relevant commits from the previous step cover 378 of 400 sampled repositories, that is, 94.5% of the repositories contain at least one README commit modifying installation-related sections. To assess whether we had missed relevant installation-related documents, for each repository, we further searched “INSTALL.md”, “INSTALLATION.md”, “docs/INSTALL.md”, and “docs/INSTALLATION.md” within its project folder (case insensitive), and consider them as additional installation documents other than the default README file. Only two repositories contain additional installation documents. Based on this observation, the majority of repositories in our analysed dataset use the README file as the primary installation documentation source, and the provided installation steps in the README file thus serve as the primary source for documentation users to access them. This observation further motivates us to focus on the analysis of the README file.
Regarding the representativeness of our analysed repositories, we selected three measures, namely repository size, total number of pull requests, and total number of issues in the repository. The repository size serves as a proxy for the complexity of the project, while the other measures quantify the project’s interaction with its users. We accessed Java and Python repositories on Github with above 2,000 stars using GHS [44]. The 2,000 stars threshold was also used in a previous study [46] when trying to build an Automatic Program Repair dataset, with the aim to collect data from popular repositories. Figure 3 shows the box plot for these three measures in our data and the repositories having above 2,000 stars. We performed the Mann-Whitney U test for these three measures, resulting in a p-value of 0.38, 0.02, and 0.06 for total size, total issues, and total pull requests, respectively. This indicates that there are no differences between the groups in the population in size and total pull requests. For total issues, we computed the Cohen’s d effect size [47], which reached -0.09, indicating that the difference is negligible. Additionally, the majority of data from both sources ranges between 100 and 1000, suggesting a medium-to-large scale of issues in both sources. This result shows that our analysed projects are comparable to popular GitHub projects in terms of their scale and user interactions.
Figure 3: Comparison between repositories in our dataset and ones above 2,000 stars
@@ III-C Dataset Characteristics
On average, each repository contains 23 relevant README commits, with a standard deviation of 25.7, indicating a skewed distribution. Specifically, with a minimum of 1 and a maximum of 224, the 25th, 50th, and 75th percentiles are 8, 14 and 31 commits, respectively.
@@
Therefore, we performed a cluster sampling [48], splitting the dataset into four clusters: repositories with README commits from the minimum to the first quartile, from the first quartile to the second quartile, from the second quartile to the third quartile, and from the third quartile to the maximum. Within each cluster, we performed a random sample with the sample size required to achieve a 95% confidence level (confidence interval: 5), resulting in samples of 206, 280, 324, and 358 README commits for each respective cluster. In total, we conducted a large-scale qualitative study on 1,168 README commits.
We calculated the correlation coefficient between the number of relevant README commits and the number of stars, as well as with the number of years the project has existed. The Pearson correlation coefficient reached 0.19 and 0.01, respectively. This statistic indicates that projects that actively participate in documentation activities, particularly for the installation-related sections, have a positive correlation to the project’s popularity, while the number of relevant README commits does not necessarily grow with the years of the project. However, it is important to note that these correlations are relatively weak, while other factors also contribute to the popularity of repositories.
@@ III-D Data Analysis
We adopted open coding [49] for qualitatively analysing the update events. The first author conducted the data analysis and created spreadsheets containing all raw data and codes. The spreadsheets were shared among all authors, and the second and third authors cross-validated all codes, concepts, and categories to reduce bias and increase the reliability of the findings [50]. Concepts and categories were discussed among authors through regular meetings and finalised after several rounds of revisions.
@@
We started our analysis by searching each commit on GitHub and recording the text summarising the changes from the perspective according to our research question: updates for the installation-related sections. We call this process “key points extraction and getting familiar with the data”. Then we rigorously performed open coding to assign each README commit with one to many succinct descriptions of their update modifications. This process took into account both the key points extracted and the content of the README commits. The additional investigation into the raw README commit data is to address any unclear key points and to double-check whether any part of the modification is missing.
During the coding procedure, we constantly compared the emerging codes within and across different README commits to synthesise a higher level of abstraction called concepts, and later the highest level of abstraction called categories. Figure 6(a) illustrates examples of the procedure of generating one of the categories called “Pre-installation Instruction Updates”. Figures 6(b) and Figure 6(c) further explain the process for how we assign codes for different changes within the commit. We have made our analysed data available. Please refer to the Replication Package section for more details.
@@
IV Results
Table II presents the hierarchical taxonomy of 189 codes that we identified from the 1,168 data points to answer our RQ1. The updates are categorised into six main groups: (1) pre-installation instruction; (2) installation-related instruction; (3) post-installation instruction; (4) help information; (5) document presentation; and (6) external resource management.
To give readers a better understanding of our analysed result, Figure 7 synthesises our analysis in a timeline fashion, corresponding to pre-installation, installation, and post-installation, where different software development tasks are included in each phase. Moreover, various information perspectives regarding the README files are also displayed within this figure, including instructions, help information, external resource links, and document presentation. For each of the categories, we describe representative examples and discuss implications for practitioners and researchers.
It is worth noting that, compared to other qualitative studies [6, 7, 1, 12], our work focuses on modifications to Markdown-formatted text, which involves two pieces of documents. Due to the substantial space to illustrate examples of README updates, we will only include figures for a few examples. Additional examples are made accessible through URLs that we have formatted as clickable links in Table II. We have also included the URLs and their corresponding screenshots in our replication package.
@@ IV Results
Category Concept # Codes Frequency Example ID Addition Modification Removal
Pre-installation instruction Dependency Setup 23 527 (45.3%) 1, 2, 3, 4 47.8% 34.8% 17.4%
Platform Setup 11 82 (7.1%) 5, 6, 7 54.5% 18.2% 27.3%
Virtual Machine and Hardware 6 19 (1.6%) —— 66.7% 16.7% 16.7%
Total 40 628 (54.0%) 52.5% 27.5% 20.0%
Installation-related instruction Source code 23 267 (23.0%) 8, 9 43.4% 43.5% 13.0%
Published package 14 162 (13.9%) 10, 11, 12, 13 57.1% 28.6% 14.3%
Total 37 429 (36.9%) 48.6% 37.8% 13.5%
Post-installation instruction Project running instruction 12 185 (15.9%) 14, 15 37.5% 50.0% 12.5%
Installation check 3 13 (1.1%) 16 40.0% 60.0% 0.0%
Running Tests 6 32 (2.8%) —— 88.9% 0.0% 11.1%
Deployment Instruction 3 9 (0.8%) —— 75% 0.0% 25.0%
Upgrade Instruction 4 6 (0.5%) —— 100.0% 0.0% 0.0%
Uninstall Instruction 2 3 (0.3%) —— 0.0% 0.0% 100.0%
Total 30 248 (21.3%) 58.6% 24.1% 17.2%
Help information Tutorial updates 6 70 (6.0%) 17 50.0% 33.3% 16.7%
Explanation 14 125 (10.7%) 18, 19 61.5% 38.5% 0.0%
Notes and troubleshooting 16 70 (6.0%) 20, 21 68.8% 6.3% 25.0%
Total 36 265 (22.8%) 62.9% 22.9% 14.3%
Document Presentation Documentation structural modification 7 197 (16.9%) 22, 23 0.0% 100.0% 0.0%
Content formatting and cleaning 10 119 (10.2%) 24, 25, 26, 27 0.0% 100.0% 0.0%
Presentation fixing 2 61 (5.2%) 28, 29 0.0% 100.0% 0.0%
Text rephrasing and editing 5 70 (6.0%) 30, 31 0.0% 100.0% 0.0%
Documentation debt 4 30 (2.6%) 32 25.0% 0.0% 75.0%
Total 28 477 (41.0%) 3.6% 85.7% 10.7%
External resources management Project external documentation 3 120 (10.3%) 33 25.0% 50.0% 25.0%
Project non-documentation artefact 11 136 (11.7%) 34 44.4% 33.3% 22.2%
Third-party external resources 4 110 (9.5%) 35 60.0% 20.0% 20.0%
Total 18 366 (31.5%) 44.4% 33.3% 22.2%
IV-A Pre-installation Instruction
@@
Dependency Setup (527). [^5] Number in the parenthesis represents the frequency of this concept. Dependencies enable code reuse and require maintenance effort as the project and other dependencies evolve [51]. In our analysed dataset, more than 40% of the documentation updates contain dependency-related instructional modifications, indicating the importance of this perspective.
@@
Apart from the programming language, other dependencies setup include external libraries and modules (e.g., API), as well as third-party software (e.g., database) that projects rely on. The inclusion of new features, the resolution of technical debt during project development, and the updates of constantly evolving dependencies collectively contribute to the volatility of the dependency management information within the README files. Indeed, most updates in this aspect involve updating dependency information, as indicated in Example 1, and instructions on installation of dependencies as indicated in Example 2. We observed a common practice of manifesting dependencies by listing them (e.g., in the forms of plain texts, bullet points, or tables), with additions, updates, and deletions to reflect the corresponding changes. Meanwhile, the patterns for providing the installation instruction of dependencies are quite diverse. Centralised commands (e.g., requirements.txt) or separate installation commands for each dependency are widely used. However, maintaining the separate command style requires more effort, as individual attention to each dependency is needed, making modifications potentially more time-consuming and the process more complicated.
In Example 3, the commit adds configuration instructions for the database on which the project relies. This underscores the importance of incorporating essential configuration details when dependencies are not ready for use following the dependency installation step, which is a common scenario for third-party software dependencies. Interestingly, some examples, such as commit Example 4, add guidelines to verify the success of dependency installation.
@@
In our analysed data, we found that 48 instances involve modification in setting up virtual environments, and the majority of them are related to newly added instructions such as the one indicated in Example 5. Furthermore, Example 6 explicitly mentions the virtual environment as their recommended way of installation of the project.
Operating system information is another platform setup concern. We observed efforts that mostly focused on adding operating system compatibility information (e.g., BiglyBT comes in several editions for different operating systems. Mac OSX, Linux and Windows use the full BiglyBT-API based on Java 8.) or updating operating system version information as indicated in Example 7, which updated the support operating system from Debian 9 and 10 to Debian 10 and 11.
@@
IV-B Installation-related Instruction
@@
Installation-related Instructions from Source Code Operations (267). Installation from source code requires users to obtain the source code and then locally execute a set of predefined operations, encompassing compilation, building, and other procedures, to install the software project. While projects may use various technologies, including Docker, Pip, and Maven, we identified in our analysed data that source code-level installation is the most commonly adopted method that the projects tend to initially provide. In addition, the inclusion of newly added instructions also accounts for various operating systems. Given that commands for Windows systems typically differ from those for Linux, additional information in the README file instructions or even the source code is needed to facilitate a smooth installation experience for different users.
Documentation efforts are also observed to provide better-encapsulated installation alternatives that are easier to install, as well as versions of projects targeting different user groups. For example, in commit Example 8, the installation option initially provided includes obtaining the source code and performing compilation. The newly added instruction introduces an alternative approach specifically for building and running a Docker image from the Dockerfile within the source code. This provides an integrated installation experience, alleviating concerns related to compilation and other dependency management issues.
Regarding update behaviours for the installation instructions from source code, a wide range of project-specific patterns are identified. For example, modifications within code instructions covering project names, file paths for executables, and configuration file names are usually trivial, reflecting changes at the file level and repository level. A simple example in commit Example 9 migrated the codebase from GitLab to GitHub, while also changing the project name from atoml-local to catlearn, leading to the corresponding adjustment of the commands in the README file. Updates to the source code installation option also involve adjustments of parameters, with various patterns observed for this type of change. As an illustrative example, commit Example 9 altered the Docker compose command with an additional “-d” command indicating a detached mode to run the Docker container.
@@
Regarding the update operations in the published package installation instructions, we identified a predominant pattern that focuses on updating the project version. Trivial updates are usually found in this type of modification, with merely a version number increment for the repository’s latest stable version, and sometimes combined with the development version. For example, both the stable version and the snapshot version were incremented within the code instruction for the Maven dependency of the repository in this commit Example 11.
Apart from version modification, other noticeable update operations for published package installation include changes for package names or package sources and changes for package management technology instructions. For example, the path to the image on DockerHub was updated in commit Example 12, reflecting the changes in the source of the project. As an example of the change in technology instruction, in commit Example 13, the Gradle command compile is updated to implementation as the previous keyword was deprecated. The rest of the updates are more specific to reflect the project development status.
Summary of installation-related instruction results: The concept of installation-related instruction from source code operation is the most commonly adopted choice in our analysed data. Their update behaviours vary, covering factors including project meta-data changes and parameter changes. On the other hand, the concept of installation-related instructions from published package operations is less frequent, and they are usually used as additional operations on top of the source code installation alternative.
IV-C Post-installation Instruction